If tbone can paint a car in 4 hours and walker texas ranger can paint the same car in 6 hours, how long will it take them to paint the car together?
Answer by stanbon(48558)

If tbone can paint a car in 4 hours and walker texas ranger can paint the same car in 6 hours, how long will it take them to paint the car together?
----------------
tbone rate is 1/4 job/hr
ranger rate is 1/6 job/hr
together rate = 1/x job/hr
-----------------------
EQUATION:
1/4 + 1/6 = 1/x
(10)/24 = 1/x
x = 24/10
x = 12/5 = 2.4 hrs/job (Time to paint the car is both work together)
=============
Cheers,
Stan H.